given the slope of the line segment and the other point, determine the unknown
(7,12), (12,d), slope 6

Answer by stanbon(75887)

given the slope of the line segment and the other point, determine the unknown
(7,12), (12,d), slope 6
----------
(d-12)/(12-7) = 6
------------
d-12 = 6(5)
------
d = 42
=================
Cheers,
Stan H.
=================